OpenStudy (anonymous):

Simplification Exponents 125^2/3

OpenStudy (koikkara):

1041.67

OpenStudy (anonymous):

its asking for fraction answer how i do that

OpenStudy (anonymous):

@Directrix

OpenStudy (er.mohd.amir):

125=5*5*5 =(5^3)^2/3 =5^2 =25

OpenStudy (anonymous):

ok whats the steps to get the answer i dont just need the answer im trying to see how to get the answer even doe its right

Directrix (directrix):

>>its asking for fraction Do you mean a common fraction or maybe a fractional exponent?

OpenStudy (anonymous):

no its not asking for one

OpenStudy (koikkara):

hmm, i get it... is it like, From this we can say that (125) ^ ⅔ is 2 times the cube root of 125. So, (5)³ * ⅔ . => (5)² => 25. So, (125) ^ ⅔ = 25

OpenStudy (koikkara):

now this looks fraction \(\huge(125) ^ ⅔\)

Directrix (directrix):

You posted this: its asking for fraction answer how i do that That is why I asked. Let me ask this. In class, are you using properties of exponents or are you converting to radicals to solve this type problem?

OpenStudy (koikkara):

then the answer is \(\huge(5)³ * ⅔ \)

OpenStudy (anonymous):

forget that i said that sorry but @Koikkara got the answer right which is 25 but i trying to figure out how to know how to get the answer

OpenStudy (koikkara):

\(\huge½\) --> square root of that number. \(\huge⅓ \)--> cube root of that number.

OpenStudy (koikkara):

125 is a cube root of 5